OpenStudy (tanvirzawad):

x[n]=x[n+r] Now, if x[n]=A*cos[w*n+B] is Discrete Time periodic signal, then prove that w*r=2*pi*q, where q is integer.

OpenStudy (anonymous):

x[n] = x[n+r] =A*cos[w*n+B] = A*cos[w*n + w*r +B] = A*cos[w*n + w*r +B] = A*cos[w*n + B + w*r] = Then use sum of angles theorem for cosine... A*cos[w*n + B]cos[w*r] - A*sin[w*n + B]sin[w*r] = Sub in wr=2*pi*q ... A*cos[w*n + B]cos[2*pi*q] - A*sin[w*n + B]sin[2*pi*q] = and evaluate ... A*cos[w*n + B]

OpenStudy (anonymous):

It's more of a "Show" than a proof, but that's how

hartnn (hartnn):

or u can simply equate the cos arguments: x[n+r] = A*cos[w*n+w*r+B] x[n]=A*cos[w*n+B] =A*cos[w*n+B+2*pi*q]<---because cos is periodic with 2pi so u have A*cos[w*n+w*r+B] = A*cos[w*n+B+2*pi*q] equating the arguments of cos, w*r = 2*pi*q

OpenStudy (anonymous):

Yeah, but also not a proof

OpenStudy (tanvirzawad):

Thank you hartnn and Tonks! May Allah Ta'ala guide you and us.